jack is 3 times as old as Lacy. 10 years from now the sum of their ages will be 60. How old are they now?

Answer by lwsshak3(11628)

jack is 3 times as old as Lacy. 10 years from now the sum of their ages will be 60. How old are they now?
let x=Lacy's age
3x=Jack's age
10 yrs from now:
x+10=Lacy's age
3x+10=Jack's age
x+10+3x+10=60
4x=40
x=10
Lacy's age=10
3x=Jack's age=30